A study to determine the validity of a new screening test for Disease X is performed on 500 study subjects. The subjects also undergo a "gold standard" diagnostic test to definitely determine the presence or absence of the disease. The prevalence of Disease X among study subjects is determined to be 10% according to the gold standard. Of the 120 subjects who test positive with the screening test, 40 are determined to have the disease according to the gold standard. Calculate the sensitivity of the screening test.

A. 80%

B. 82%

C. 33%

D. 97%

E. 68%
Answer #1

Ans.) 50%

Solutions:

Sensitivity of tests = True positive / (True positive + False negative)

True positives are the ones which give positive result for given test and are also actually positive for the given condition.

False negative are the ones which give negative result for the tests but are actually positive for the given condition.

False positive are the ones that give positive test result even if they are actually negative for the given condition

True negative are the ones that give negative result and also are actually negative for the given condition

According to standard gold test, total number of actual postive test = 10% of 500 = 50

According to conducted new test:

Number of true positives = 40

Number of false negative = 10

So, sensitivity of new test = 10 / (10+40) = 10/50 = 50%